1 The Romanian language

The file romanian.dtx^{1} defines all the language-specific macros for the Romanian language.

For this language currently no special definitions are needed or available.

The macro \LdfInit takes care of preventing that this file is loaded more than once, checking the category code of the @ sign, etc.

1 ⟨∗code⟩

 2 \LdfInit{romanian}\captionsromanian

When this file is read as an option, i.e. by the \usepackage command, romanian will be an ‘unknown’ language in which case we have to make it known. So we check for the existence of \l@romanian to see whether we have to do something here.

 3 \ifx\l@romanian\@undefined
 4  \@nopatterns{Romanian}
 5  \adddialect\l@romanian0\fi

The next step consists of defining commands to switch to (and from) the Romanian language.

\captionsromanian The macro \captionsromanian defines all strings used in the four standard documentclasses provided with L^{A}TEX.

 6 \addto\captionsromanian{%
 7 \def\prefacename{Prefa\c{t}\u{a}}%
 8 \def\refname{Bibliografie}%
 9 \def\abstractname{Rezumat}%
10 \def\bibname{Bibliografie}%
11 \def\chaptername{Capitolul}%
12 \def\appendixname{Anexa}%
13 \def\contentsname{Cuprins}%
14 \def\listfigurename{List\u{a} de figuri}%
15 \def\listtablename{List\u{a} de tabele}%
16 \def\indexname{Glosar}%
17 \def\figurename{Figura}%    % sau Plan\c{s}a
18 \def\tablename{Tabela}%
19 \def\partname{Partea}%
20 \def\enclname{Anex\u{a}}%   % sau Anexe
21 \def\ccname{Copie}%
22 \def\headtoname{Pentru}%
23 \def\pagename{Pagina}%
24 \def\seename{Vezi}%
25 \def\alsoname{Vezi de asemenea}%
26 \def\proofname{Demonstra\c{t}ie} %
27 }%

\dateromanian The macro \dateromanian redefines the command \today to produce Romanian dates.

28 \def\dateromanian{%
29 \def\today{\number\day~\ifcase\month\or
30 ianuarie\or februarie\or martie\or aprilie\or mai\or

1The file described in this section has version number v1.2h and was last revised on 1996/12/23. A contribution was made by Umstatter Horst (hhu@cernvm.cern.ch).

31 iunie\or iulie\or august\or septembrie\or octombrie\or
32 noiembrie\or decembrie\fi
33 \space \number\year}}

\extrasromanian \noextrasromanian The macro \extrasromanian will perform all the extra definitions needed for the Romanian language. The macro \noextrasromanian is used to cancel the actions of \extrasromanian For the moment these macros are empty but they are defined for compatibility with the other language definition files.

34 \addto\extrasromanian{}
35 \addto\noextrasromanian{}

The macro \ldf@finish takes care of looking for a configuration file, setting the main language to be switched on at \begin{document} and resetting the category code of @ to its original value.

36 \ldf@finish{romanian}

37 ⟨/code⟩